### v1

# Deterministic Discrete Denoising

Abstract
We propose a deterministic denoising algorithm for discrete-state diffusion models based on Markov chains. The generative reverse process is derandomized by introducing a variant of the herding algorithm with weakly chaotic dynamics, which induces deterministic discrete state transitions. Our approach is a direct replacement for the stochastic denoising process, requiring neither retraining nor continuous state embeddings. We demonstrate consistent improvements in both efficiency and sample quality on text and image generation tasks. Thus, this simple derandomization approach is expected to enhance the significance of discrete diffusion in generative modeling. Furthermore, our results reveal that deterministic reverse processes, well established in continuous diffusion, can also be effective in discrete state spaces.
1 Introduction
Diffusion models have recently achieved remarkable success in generating realistic image and audio data. Their forward process progressively corrupts data with noise, and the models are trained to denoise by inverting the corruption process. Once trained, diffusion models can generate realistic samples from pure noise by iteratively applying the denoising procedure. For continuous data such as images and speech, the reverse denoising process is often implemented as a deterministic algorithm, which exhibits more directed dynamics and enables efficient generation with fewer denoising steps compared to stochastic denoising, as demonstrated by DDIM (Song et al., 2021).
Diffusion models for discrete data are also an important research direction, since many data of interest are inherently discrete, including text, graphs, genomes, and chemical structures. In discrete-state diffusion models, the forward and reverse processes are typically formulated as probabilistic Markov chains on a discrete sample space (Austin et al., 2021). If the reverse denoising process could be made deterministic, improved performance would be expected, as in the case of continuous models. However, in discrete spaces, naive derandomization as a deterministic mapping aggregates discrete trajectories and reduces sample diversity, which is essential for generative modeling.
One promising approach is to embed the discrete sample space into a continuous space and apply continuous diffusion models (Hoogeboom et al., ; Chen et al., 2023; Sahoo et al., 2025). The main advantage of this approach is that it allows the direct use of well-established techniques for continuous diffusion, including deterministic denoising. Nevertheless, denoising in the continuous space requires retraining the model. Moreover, it remains unclear whether discrete data must be embedded into a continuous space for derandomization.
In this paper, we propose to directly derandomize the reverse denoising process on discrete sample spaces. The key idea is to introduce a variant of the herding algorithm (Welling, 2009; Welling & Chen, 2010), which associates auxiliary continuous weight variables with each discrete variable in the sample space. The reverse denoising updates of both discrete and continuous variables are deterministic, and randomness arises solely from their initial states. Since the proposed algorithm derandomizes the state transitions of Markov chains, it serves as a drop-in replacement for the stochastic reverse process without retraining discrete diffusion models. This relation is analogous to DDIM (Song et al., 2021), which generates samples deterministically using a model trained for DDPM (Ho et al., 2020). In our experiments on text and image generation, we demonstrate consistent improvements in both efficiency and sample quality. These results reveal that deterministic reverse processes, well established in continuous diffusion, can also be effective in discrete state spaces.
In the evaluation, we adopt UDLM (Schiff et al., 2025), a state-of-the-art discrete-state generative model based on the uniform diffusion process. Our aim in this paper is not to propose a new discrete diffusion model, but to demonstrate how the herding-based algorithm improves the generative process of existing ones. We use the official training, sampling, and evaluation codes of UDLM, replacing only the reverse denoising procedure. This directly shows that our method works as a drop-in replacement without retraining. Moreover, the uniform diffusion process is particularly suitable for our approach, since its reverse process performs self-correction through successive Markov state transitions, making it a natural setting to examine whether derandomizing these transitions can improve efficiency and sample quality.
In contrast, masked diffusion has been widely studied as a discrete diffusion model, but its reverse process allows each token to transition only once from the masked state and remain fixed thereafter, fundamentally limiting its dynamics. A related line of work is the deterministic reverse algorithm proposed in Chen et al. (2024), which performs deterministic state transitions in discrete sample spaces. This approach is effective for masked diffusion, where the transition time is limited. Our work complements this result by showing that deterministic denoising is also beneficial for uniform diffusion models, where state transitions continue throughout the reverse process.
Overall, we present a simple yet effective derandomization approach to directly enhance the generative process of existing discrete-state diffusion models, thus strengthening the significance of discrete diffusion in generative modeling.
2 Background
2.1 Discrete Diffusion Models
We consider generative models for discrete data represented as a sequence of $L$ categorical tokens. Each token takes one of $K$ categories and is represented as a $K$-dimensional one-hot vector $\bm{x}\in\mathcal{V}$, where $\mathcal{V}=\{(x_{1},x_{2},\dots,x_{K})^{\top}\in\{0,1\}^{K}\mid\sum_{k=1}^{K}x_{k}=1\}$. A sequence of $L$ tokens is denoted by $\bm{x}^{(1:L)}=(\bm{x}^{(1)},\bm{x}^{(2)},\dots,\bm{x}^{(L)})\in\mathcal{V}^{L}$.
The forward noising process in the D3PM framework (Austin et al., 2021) is formulated as a discrete-time Markov chain on $\mathcal{V}$, which gradually transforms the original data $\bm{x}_{0}$ into noise $\bm{x}_{T}$. Given the Markov transition matrix $Q_{t}$ at time $t$, the forward transition is defined as
$q(\bm{x}_{t}\mid\bm{x}_{t-1})=\mathrm{Cat}(\bm{x}_{t};Q_{t}\bm{x}_{t-1}),$ (1)
where $\mathrm{Cat}(\cdot;\bm{p})$ denotes the categorical distribution with the probability vector $\bm{p}$.
The Markov transition matrix $Q_{t}$ is designed to gradually transform any categorical distribution into a stationary distribution as $t$ increases. For the uniform noising process, it is given by
$Q_{t}=(1-\beta_{t})\bm{I}+\beta_{t}\frac{1}{K}\bm{1}\bm{1}^{\top},$ (2)
where $\beta_{t}\in(0,1)$ is a noise schedule. For the absorbing noising process, it is defined as
$Q_{t}=(1-\beta_{t})\bm{I}+\beta_{t}\bm{x}_{\mathrm{mask}}\bm{1}^{\top},$ (3)
where $\bm{x}_{\mathrm{mask}}$ is the one-hot vector representing the absorbing category. In both cases, the Markov transition matrix from time $s$ to $t$, $Q_{t}Q_{t-1}\dots Q_{s+1}$, admits a closed-form expression for arbitrary $s<t$, which can be extended to non-integer times. This property is exploited to improve the efficiency of training and sampling (Sahoo et al., 2024; Schiff et al., 2025).
The reverse denoising process is also formulated as a discrete-time Markov chain on $\mathcal{V}^{L}$. A neural network model parameterized by $\theta$ is trained to predict $\tilde{\bm{x}}^{(1:L)}_{0}$, an estimate of the original data $\bm{x}^{(1:L)}_{0}$, from noisy data $\bm{x}^{(1:L)}_{t}$ at time $t$. Training is performed by minimizing a variational upper bound (NELBO) on the negative log-likelihood (Austin et al., 2021; Schiff et al., 2025). The transition probability $p_{\theta}(\bm{x}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$ of the reverse process is then computed from the predicted original data $\tilde{\bm{x}}^{(1:L)}_{0}$ and the forward transition matrix $Q_{t}$.
By iterating this Markov transition from pure noise $\bm{x}^{(1:L)}_{T}$ sampled from the stationary distribution, the reverse process generates a sample $\bm{x}^{(1:L)}_{0}$. Namely, each denoising step reduces to sampling from the categorical distribution $\mathrm{Cat}(\bm{x}_{t-1};\bm{p}_{t-1})$ with probability vector $\bm{p}_{t-1}=p_{\theta}(\bm{x}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$. This is typically implemented as a probabilistic procedure using the Gumbel-max trick. However, sampling does not necessarily have to be probabilistic; it suffices that samples are generated faithfully according to the given transition probabilities. This motivates us to explore deterministic alternatives for the reverse denoising process.
2.2 Herding Algorithm
The herding algorithm (Welling, 2009; Welling & Chen, 2010) is a deterministic sampling method with weakly chaotic dynamics, which combines statistical learning and inference for energy-based models into a single dynamical system.
The herding system yields a sample sequence $\bm{x}_{1},\bm{x}_{2},\dots$ from a discrete sample space $\mathcal{V}$ such that the empirical average of feature values $\bm{\phi}(\bm{x}_{t})$ converges to predefined target values $\bm{\mu}\in\mathbb{R}^{N}$, where $\bm{\phi}\colon\mathcal{V}\to\mathbb{R}^{N}$ is a set of real-valued feature functions.
To reduce the discrepancy between the averaged feature values and the target values in each sampling step, the herding system updates the weight vector $\bm{w}_{t}$. Specifically, the herding system generates a new sample $\bm{x}_{t+1}$ and updates the weight vector $\bm{w}_{t+1}$ at time $t+1$ as follows:
$\bm{x}_{t+1}=\operatorname*{arg\,max}_{\bm{x}\in\mathcal{V}}\bm{w}_{t}^{\top}\bm{\phi}(\bm{x}),$ (4)
$\bm{w}_{t+1}=\bm{w}_{t}+\bm{\mu}-\bm{\phi}(\bm{x}_{t+1}).$ (5)
This forms a hybrid dynamical system (Aihara & Suzuki, 2010) consisting of a discrete state variable $\bm{x}_{t}$ and a continuous state variable $\bm{w}_{t}$. Given an initial state $(\bm{x}_{0},\bm{w}_{0})\in\mathcal{V}\times\mathbb{R}^{N}$, the system generates samples iteratively, deterministically, and autonomously. Typically, the weight vector $\bm{w}_{t}$ does not converge and remains in a bounded domain around the origin.
The difference between the predefined target vector and the feature vector averaged over $T$ samples can be written as
$\bm{\mu}-\frac{1}{T}\sum_{t=1}^{T}\bm{\phi}(\bm{x}_{t})=\frac{1}{T}(\bm{w}_{T}-\bm{w}_{0}).$ (6)
The right-hand side converges to zero as $T\to\infty$, provided that the sequence $\bm{w}_{t}$ remains bounded, which is guaranteed under mild conditions. Therefore, the herding system generates a sequence of samples with the predefined feature expectations converging asymptotically at a rate of $O(T^{-1})$. This convergence rate is much faster than the $O(T^{-1/2})$ rate of probabilistic sampling, such as independent categorical sampling and Markov chain Monte Carlo (MCMC) sampling, from a distribution with the same feature expectations. Note that exact maximization in the sample generation step is not required, since the $O(T^{-1})$ convergence holds as long as the weight vector remains bounded (Welling, 2009; Welling & Chen, 2010).
The dynamics of the herding system is classified as a piecewise isometry (Goetz, 2000; 2003), which is weakly chaotic, with Lyapunov exponents equal to zero almost everywhere, and typically has a fractal attracting set. Such complex nonlinear dynamics of the herding system yields diverse samples with relatively high entropy and negative autocorrelations (Welling, 2009; Welling & Chen, 2010).
For sampling from a categorical distribution $\mathrm{Cat}(\bm{x};\bm{p})$ on $\mathcal{V}$, the feature function $\bm{\phi}(\bm{x})=\bm{x}$ and the target vector $\bm{\mu}=\bm{p}$ lead to the following herding dynamics:
$\bm{x}_{t+1}=\operatorname*{arg\,max}_{\bm{x}\in\mathcal{V}}\bm{w}_{t}^{\top}\bm{x},$ (7)
$\bm{w}_{t+1}=\bm{w}_{t}+\bm{p}-\bm{x}_{t+1}.$ (8)
This system generates samples from $\mathrm{Cat}(\bm{x};\bm{p})$ whose empirical distribution is closer to $\bm{p}$ than that of probabilistic categorical sampling.
3 Derandomization of Reverse Denoising Process
We apply the herding algorithm to the reverse denoising process of discrete diffusion models. For derandomization of the Markov chain, we need to introduce a time-dependent version of the herding algorithm as in (Suzuki, 2014).
Specifically, the sample and weight vector are updated from time $t$ to t1t-1 as follows:
$\bm{x}_{t-1}$ $=\operatorname*{arg\,max}_{\bm{x}\in\mathcal{V}}(\bm{w}_{t}+\bm{p}_{t-1})^{\top}\bm{x},$ (9)
$\bm{w}_{t-1}$ $=\bm{w}_{t}+\bm{p}_{t-1}-\bm{x}_{t-1},$ (10)
where $\bm{p}_{t-1}=p_{\theta}(\bm{x}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$ denotes the probability vector of the categorical distribution obtained from the trained model’s prediction. This herding-based system yields a sample sequence that more faithfully reflects transition probabilities than probabilistic sampling, reducing the discrepancy $\sum_{\tau=t}^{T-1}(\bm{p}_{\tau}-\bm{x}_{\tau})$.
To mitigate chattering of the discrete states, we introduce a delayed-switching mechanism with a margin $\delta>0$: the sample $\bm{x}_{t-1}$ is kept equal to $\bm{x}_{t}$ unless another candidate in the argmax step exceeds the objective value for $\bm{x}_{t}$ by at least $\delta$.
The algorithm is entirely deterministic; the only source of randomness arises from the initialization of $(\bm{x}_{T},\bm{w}_{T})$, which is chosen randomly from the uniform distribution on $\mathcal{V}\times[0,1]^{K}$. By iterating the update rules, the algorithm induces a mapping from a pure noise $(\bm{x}^{(1:L)}_{T},\bm{w}^{(1:L)}_{T})$ to a sample $(\bm{x}^{(1:L)}_{0},\bm{w}^{(1:L)}_{0})$, which constitutes a piecewise isometry. Unlike the mapping defined by ODE flows in continuous diffusion models, it is not one-to-one and the probability mass is preserved piecewise. In other words, small perturbations in the weight vector persist unless they alter the argmax; however, once the generated sample changes, they cause a drastic shift in the trajectory of the dynamics, reflecting the weakly chaotic behavior with zero Lyapunov exponents almost everywhere.
The time-dependent herding system introduced here can be regarded as a natural extension of the $\Delta\Sigma$-modulator (Inose et al., 1962; Inose & Yasuda, 1963) and the error diffusion algorithm (Adler et al., 2003), which have long been studied in engineering fields such as signal and image processing.
4 Experiments
We demonstrate the effectiveness of the proposed discrete denoising algorithm through experiments on text and image generation. For clarity of presentation, we reproduce the trained UDLM model as reported in (Schiff et al., 2025) and evaluate the improvements that result solely from introducing the herding algorithm into the reverse denoising process. All experiments are conducted based on the official code provided by Schiff et al. (2025).
4.1 Character-Level Text Generation
We first evaluate the proposed algorithm in character-level text generation, using the text8 dataset (Mahoney, 2011). The vocabulary size is $35$ and the output length is fixed to $256$ characters. The model architecture follows the UDLM paper (Schiff et al., 2025), which is based on Transformer with 92.4M parameters. For evaluation, we generate 1,024 samples for each condition and compute the generative perplexity based on the GPT-2 Large pretrained model (Radford et al., 2019) and the entropy.
Figure 1 shows the generative perplexity and entropy of the generated samples. The herding-based denoising with $\delta=0$ consistently outperforms stochastic denoising in terms of perplexity, achieving up to a $10\times$ improvement for $T=1024$ inference steps, albeit at the expense of entropy. Increasing the margin $\delta$ interpolates between these two extremes, allowing us to improve perplexity with a moderate trade-off in entropy.
Figure 1: Evaluation of text8 character-level text samples by generative perplexity (Gen. PPL) and entropy. Results for stochastic denoising (Gumbel-max) and herding-based denoising with different $\delta$ values are shown.
4.2 Word-Level Text Generation
We also evaluate the proposed algorithm in word-level text generation, using the one billion words dataset (LM1B) (Chelba et al., 2013). The vocabulary size is $30,522$ and the output length is $128$ tokens. The model architecture follows the UDLM paper (Schiff et al., 2025), which is based on Transformer with 139M parameters. For evaluation, we generate 1,024 samples for each condition and compute the generative perplexity based on the GPT-2 Large pretrained model (Radford et al., 2019) and the entropy.
Table 1 shows the evaluation results. Our experiments are based on the HuggingFace (HF) UDLM checkpoint trained with LM1B provided by Schiff et al. (2025). We note that the HF model appears to differ from the one used in their PPL evaluation, as the generative perplexity is substantially worse than the published values. The herding-based denoising significantly improves the perplexity compared to stochastic denoising for both $T=1024$ and $T=128$ steps, and even outperforms the published UDLM result for $T=1024$ steps, closing the gap to the autoregressive (AR) model. We also observe a slight trade-off between the perplexity and entropy, as reported in the OWT experiment of Duo (Sahoo et al., 2025).
Table 1: Evaluation of LM1B text samples by generative perplexity (Gen. PPL) and entropy. The bottom two rows show our results using the HF trained model from Schiff et al. (2025). The herding margin is set to $\delta=0.0005$ for $T=1024$ and $\delta=0.01$ for $T=128$. Best Gen. PPL values (excluding AR) are indicated in bold. Published values from Schiff et al. (2025).
$T=1024$ $T=128$
Method Gen. PPL ($\downarrow$) Entropy ($\uparrow$) Gen. PPL ($\downarrow$) Entropy ($\uparrow$)
AR 067.46
MDLM 116.80 120.93
UDLM 078.22 079.87
UDLM (HF model) 180.15 6.88 184.54 6.87
UDLM herding (ours) 074.95 6.31 108.25 6.40
4.3 Categorical-Valued Image Generation
We apply our herding-based algorithm to image generation, following the experiments of D3PM (Austin et al., 2021) and UDLM (Schiff et al., 2025). While pixel values are ordinal by nature, they are treated as discrete categorical variables for the purpose of evaluating generative performance. Namely, the forward uniform noising process assigns equal probability to transitions to neighboring and distant pixel values alike.
The dataset is discretized CIFAR-10 (Krizhevsky & Hinton, 2009), where the pixel intensities of the $32\times 32$ RGB images are discretized to 8-bit integer values. The model is based on a U-Net architecture (Ronneberger et al., 2015), which is identical to that used in Schiff et al. (2025). For evaluation, we computed Fréchet inception distance (FID) (Heusel et al., 2017) and inception score (IS) (Salimans et al., 2016) on 50,000 generated samples, using the PyTorch evaluation code  [^1] Available at https://github.com/w86763777/pytorch-image-generation-metrics with a pretrained Inception v3 model (Szegedy et al., 2016).
From the evaluation results shown in Table 2, we confirm that the trained unconditional UDLM model and conditional UDLM model (D-CFG with $\gamma=1$) reproduce the published values nearly identically. The proposed herding-based denoising consistently outperforms stochastic denoising in both FID and IS across all cases. Remarkably, the herding-based denoising with $T=128$ steps achieves performance comparable to that of stochastic denoising with $T=1024$ steps.
Table 2: Evaluation of CIFAR-10 image samples by FID and IS. The bottom two rows show our results using the UDLM models trained with the code from Schiff et al. (2025). D-CFG denotes conditional models with $\gamma=1$. The herding margin is set to $\delta=0.07$ for $T=128$ and $\delta=0.15$ for the other cases. Best values are indicated in bold. Published values from Austin et al. (2021) and Schiff et al. (2025).
$T=1000$ $T=1024$ (D-CFG) $T=128$ (D-CFG)
Method FID ($\downarrow$) IS ($\uparrow$) FID ($\downarrow$) IS ($\uparrow$) FID ($\downarrow$) IS ($\uparrow$)
D3PM Uniform 51.27 5.99
MDLM 33.75 6.74 27.94 7.14 64.09 5.81
UDLM 33.65 6.86 26.70 7.43 30.48 7.30
UDLM (reproduced) 32.64 6.97 24.94 7.41 29.19 7.18
UDLM herding (ours) 26.39 7.40 19.20 7.79 23.95 7.51
5 Discussion
The deterministic denoising algorithm defines a mapping from the initial states $(\bm{x}^{(1:L)}_{T},\bm{w}^{(1:L)}_{T})$ to their final states $(\bm{x}^{(1:L)}_{0},\bm{w}^{(1:L)}_{0})$, which induces a flow of probability mass within the discrete state space $\mathcal{V}^{L}$. Each initial state encodes a generated sample $\bm{x}^{(1:L)}_{0}$, whose required precision is determined by the granularity of the piecewise isometric mapping.
The initial weight vectors $\bm{w}_{T}$ are currently chosen randomly from the uniform distribution on $0,1]^{K}$. Scaling the initial weights can influence the balance between the determinism and randomness of the algorithm. Moreover, there is no reason that all time steps should be treated equally in the accumulation of the discrepancy between the empirical and the target distributions. It is natural to consider weighted accumulation, where the weight scheduling can be an important factor in denoising performance.
We found that introducing delayed switching in the herding algorithm is effective in improving the quality of generated samples. However, it is not clear how to set the appropriate switching margin $\delta$, which can depend on the model, the dataset, the vocabulary size $K$, and the number of inference steps $T$. Alternative mechanisms for delayed switching are also possible. Further theoretical and numerical studies will be necessary to clarify the role of delayed switching in deterministic denoising.
These observations on weighted accumulation and delayed switching naturally lead to the idea of continuous-time formulation. Discrete diffusion models formulated as continuous-time Markov chains (CTMCs) have been studied extensively (Campbell et al., 2022; Lou et al., 2024), and can also be naturally derandomized using the continuous-time variant of the herding system, which exhibits chaotic billiard dynamics (Suzuki, 2015). As an alternative, chaotic billiard sampling (Suzuki et al., 2013; Lee et al., 2025), which can be extended to multi-state spin models such as the Potts model (Suzuki, 2013), can also be applied. Compared with the CTMC formulation, these derandomized billiard systems require computing the exact state transition times induced by collisions with the billiard boundary, which is more time-consuming. Since the exact computation may not contribute to the performance, time-discretized computation is considered more practical. Nonetheless, continuous-time formulations may offer advantages in dedicated hardware implementations, as demonstrated in the integrated-circuit implementation of chaotic Boltzmann machines (Yamaguchi et al., 2019).
As mentioned in the introduction, we focused on the uniform diffusion process in this paper. However, the proposed herding-based denoising algorithm can also be applied to other Markov-chain formulations of discrete diffusion models. While in Chen et al. (2024) the state transition time from the absorbing state is externally specified, the proposed deterministic algorithm may allow the transition time to be determined by the system itself according to the transition probability.
The capability of the herding algorithm is not limited to sampling from categorical distributions. By appropriately designing feature functions, additional constraints, such as symmetry, can be incorporated into the generated samples, which may further enhance the applicability of discrete diffusion models. Although we presented a fully deterministic algorithm, it is also possible to combine the method with probabilistic sampling. Since the boundedness of the weight vector is essential, monitoring its dynamics may help improve the performance of stochastic denoising, as in bounded-error sampling (Yamashita & Suzuki, 2019).
In summary, we presented a novel strategy for derandomizing the reverse denoising process of discrete diffusion models, based on the herding algorithm with weakly chaotic dynamics. Our results provide a positive answer to the key question of whether deterministic denoising can also be effective in discrete diffusion models. As our proposed algorithm is minimal, there are many directions for further improvement. We expect that our approach will contribute to enhancing the significance of discrete diffusion in generative modeling.
Acknowledgments
This work was supported by ALCA-Next JPMJAN23F2 and Moonshot R&D JPMJMS2021 from the Japan Science and Technology Agency (JST), by Project JPNP14004 from the New Energy and Industrial Technology Development Organization (NEDO), and by KAKENHI 25K21300 from Japan Society for the Promotion of Science (JSPS).